# Pseudoentanglement from tensor networks

### v2

@@ Abstract
Introduction.—Entanglement is a fundamental resource in quantum information science and an important principle in theoretical physics, from condensed matter to gravity Horodecki and Horodecki (2009); Kitaev and Preskill (2006); Levin and Wen (2006); Swingle (2012); Maldacena and Susskind (2013); Liu and Suh (2014); Kaufman et al. (2016); Abanin et al. (2019). At the same time, the entanglement structure of large many-body states is very challenging to observe. Specifically, there exist $N$-qubit states with very low entanglement that are computationally indistinguishable (i.e., indistiguishable to any observer whose resources scale polynomially in $N$) from states with much higher entanglement, such as Haar-random states. This property is knowns as pseudoentanglement Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Bouland et al. (2023).
@@
Motivated originally by ideas in quantum cryptography Ji et al. (2018), the concept of pseudoentanglement enables connections to many-body physics in contexts including thermal equilibrium in many-body dynamics Feng and Ippoliti (2024), learning of Hamiltonian ground states Bouland et al. (2023), and holography Aaronson et al. (2023), where it may have implications for the complexity of the AdS/CFT “dictionary” Bouland et al. (2019). In general, it is a powerful tool to demonstrate the existence of entanglement structures that are computationally inaccessible, and thus contributes to sharpen the limits of what is learnable in realistic quantum experiments.
@@
Existing constructions of pseudoentangled state ensembles are based on subset-phase-states, |$\ket{\psi_{\mathcal{S},f}}\propto\sum_{x\in\mathcal{S}}(-1)^{f(x)}\ket{x}$, where under suitable constraints on the Boolean function $f:\{0,1\}^{N}\to\{0,1\}$ and/or the subset $\mathcal{S}\subseteq\{0,1\}^{N}$ one can prove computational pseudorandomness and enforce an upper bound on entanglement either on specific cuts or on all cuts at once Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023). These constructions are very powerful and in some cases achieve the theoretical maximum amount of hideable entanglement. However, they natively represent unstructured states where all $N$ qubits are on the same footing, without e.g. a notion of locality. This makes them well suited to emulate maximally random states such as those that describe infinite-temperature equilibrium Feng and Ippoliti (2024), but limits their ability to mimic more structured systems. While it is possible to tailor phase-states to reflect some locality structure Bouland et al. (2023), it remains not clear in general how to mimic physical systems of interest, such as holographic ones.
@@
Motivated by these considerations, in this work we introduce new families of pseudoentangled states with a highly flexible entanglement structure. We leverage tensor networks Cirac et al. (2021); Evenbly and Vidal (2011); Zaletel and Pollmann (2020); Jahn and Eisert (2021); Haghshenas et al. (2022), which allow us to hard-code locality and the associated entanglement (pseudo-)area-law into our states. Recently proposed efficient constructions of pseudorandom unitaries Metger et al. (2024) then allow for the efficient preparation of pseudoentangled states whose true (computationally hidden) entanglement structure is that of an arbitrary tensor network that admits an isometric realization Zaletel and Pollmann (2020); Haghshenas et al. (2022); Soejima et al. (2020); Tepaske and Luitz (2021). These include many examples of interest, from matrix product states (MPS) to projected entangled pair states (PEPS) in any dimension Cirac et al. (2021); Verstraete et al. (2006); Schuch et al. (2007) and even holographic tensor network states Pastawski et al. (2015); Hayden et al. (2016); Jahn and Eisert (2021); Vasseur et al. (2019).
@@
in terms of the ‘moment operators’ $\rho^{(m)}_{\mathcal{E}}=\mathbb{E}_{\psi\sim\mathcal{E}}[\outerproduct{\psi}{\psi}^{\otimes m}]$. This means that an observer with access to polynomially many copies $m$ of a given state would have no way of deciding which ensemble it came from in polynomial time. The constraint on entanglement is simply that, for some extensive subsystem $A$ and with high probability over the state $\psi\sim\mathcal{E}$, the entanglement entropy $S(A)$ is bounded above by a function $f(N)\leq o(N)$. By comparison, the scaling of entanglement entropy in Haar-random states (known as the Page curve Page (1993)) is $S(A)=\Theta(N)$ for any extensive subsystem. The discrepancy between true and apparent entanglement is called the pseudoentanglement gap. The theoretically maximal pseudoentanglement gap is $f(N)=\omega(\log N)$ vs $\Theta(N)$, and can be saturated by random subset-phase states Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Bouland et al. (2023).
@@
Pseudoentangled matrix product states.—We first exhibit our construction on the simplest tensor network: the one-dimensional matrix product state (MPS). A MPS, sketched in Fig. 1(a), is a many-body state whose amplitudes are expressed as $\innerproduct{i_{1},\dots i_{N}}{\psi}={\rm Tr}(A^{[1]}_{i_{1}}\cdots A^{[N]}_{i_{N}})$, with each $A^{[j]}_{i_{j}}$ a $\chi\times\chi$ matrix; $\chi$ is a numerical cut-off known as the bond dimension, which upper-bounds the Schmidt rank of the state about any bond on the 1D line and thus imposes an area-law, Fig. 1(b). A standard route to prepare MPSs in the lab is to use ‘staircase’ circuits  [^2] This approach is also known as “holographic” foss-feig_holographic_2021; anand_holographic_2023, though not to be confused with the other usage of the word in this paper., where one associates $A_{i,\alpha,\beta}^{[j]}=\bra{i,\beta}U^{[j]}\ket{0,\alpha}$, with $U^{[j]}$ a unitary gate acting on a qubit and a $\chi$-state qudit. This is sketched in Fig. 1(c). In this work we take $\chi=2^{\nu}$, so each $U^{[j]}$ is a $(\nu+1)$-qubit gate. The staircase circuit comprises $L$ gates; for the last gate, all $\nu+1$ output qubits become part of the physical system, which in all is made of $N=L+\nu$ qubits. We take our unitary gates $\{U^{[j]}\}_{j=1}^{L}$ to be drawn from the recently introduced ‘PFC ensemble’ Metger et al. (2024) of pseudorandom unitaries (PRUs). These are efficiently implementable transformations that are computationally indistinguishable from Haar-random unitaries (see  [^3] See Supplementary Information., Metger et al. (2024) for more details). We denote the resulting ensemble of $N$-qubit pure states by $\mathcal{E}_{\rm pMPS}$.
@@
If $\omega({\rm poly}(N))\leq\chi\leq 2^{o(N)}$, the pseudorandom MPS ensemble $\mathcal{E}_{\rm pMPS}$ is pseudoentangled.
@@
We split this task in three parts. First, by definition of the PFC ensemble Metger et al. (2024), the pseudorandom ensemble $\mathcal{E}_{\rm pMPS}$ is computationally indistinguishable from its counterpart $\mathcal{E}_{\rm rMPS^{\prime}}$ where the pseudorandom PFC gates are replaced by genuinely random permutations, phase, and Clifford gates. Next, we prove that $\mathcal{E}_{\rm rMPS^{\prime}}$ is indistinguishable from its counterpart made of genuine Haar-random gates, which we term $\mathcal{E}_{\rm rMPS}$:
@@
We have $\|\rho^{(m)}_{\mathcal{E}_{\rm rMPS^{\prime}}}-\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}\|_{\rm tr}\leq O(Nm/\sqrt{\chi})$.
The proof of this fact follows easily from the properties of the PFC ensemble, and is reported in Note (2). We finally show that the random MPS ensemble $\mathcal{E}_{\rm rMPS}$ is indistinguishable from the Haar-random ensemble:
@@
    the $m$-th twirling channel for the Haar measure is given by $\Phi^{(m)}_{\rm Haar}(O)=\mathbb{E}_{U\sim{\rm Haar}}[U^{\otimes m}O(U^{\dagger})^{\otimes m}]=\sum_{\sigma,\tau\in S_{m}}\textsf{Wg}_{q}(\sigma\tau^{-1}){\rm Tr}(O\hat{\tau})\hat{\sigma}$, where $\textsf{Wg}_{q}(\sigma\tau^{-1})$ is the Weingarten function, see Note (2).
@@
where $D=2^{N}$ is the physical Hilbert space dimension, each $\hat{\sigma}$ denotes a replica permutation for a single qubit ($i<L$) or $2\chi$-state qudit ($i=L$), and $e^{-E[\{\sigma_{i},\tau_{i}\}]}$ is a “Boltzmann weight” (not necessarily positive) of a configuration of $S_{m}$-valued spins on a line, sketched in Fig. 1(d). This way of writing the operator is reminiscent of a partition function Hunter-Jones (2019) and will be helpful in our analysis.
@@
We first show that $\Delta_{m}^{\rm u.}$ is small. The Boltzmann weight of the uniform spin configuration is $e^{-E[\{\sigma,\sigma\}]}=\chi^{L-1}\textsf{Wg}_{2\chi}(e)^{L}$ (independent of $\sigma$), with $e$ the identity permutation. We show in Note (2) that $\textsf{Wg}_{2\chi}(e)=(2\chi)^{-m}[1+O(m^{2}/\chi)]$; therefore $e^{-E[\{\sigma,\sigma\}]}=D^{-m}[1+O(m^{2}/\chi)]$. At the same time, by the ‘birthday asymptotics’ we have $f_{D,m}=D^{-m}[1+O(m^{2}/\chi)]$, so overall
@@
Holographic tensor networks.—An important application of our construction is to tensor network models of holography, where a quantum system is viewed as the “boundary” of a higher-dimensional “bulk” space, and its entanglement structure reflects the geometry of the bulk. These models Pastawski et al. (2015); Swingle (2012); Qi (2013); Hayden et al. (2016); Qi et al. (2017); Jahn and Eisert (2021) have attracted considerable interest as toy models of the AdS/CFT correspondence and its links to quantum error correction. They generally feature tensors arranged on a tiling of $2D$ hyperbolic space, which may or may not have physical legs in the bulk, terminating with “dangling” legs on the boundary which represents the physical system of interest. Holographic tensor networks are very well suited to our construction, since they encode the bulk isometrically into the boundary; i.e., the tensor network may be viewed as a circuit propagating information unitarily outward from the center of the hyperbolic space toward its boundary. (We do not consider recently introduced non-isometric codes here, though that would be an interesting direction for future work Akers et al. (2022); DeWolfe and Higginbotham (2023).)
@@
Here we focus for concreteness on the $\{6,4\}$ tiling of the hyperbolic plane with 6-leg tensors (3-qudit unitary gates), without bulk degrees of freedom. This is sketched in Fig. 2(b). Placing a PRU of dimension $\chi^{3}$ on each hexagon defines a state ensemble $\mathcal{E}_{\rm holo}$. Our general proof extends to this ensemble and is sufficient to show pseudoentanglement: $\mathcal{E}_{\rm holo}$ is computationally indistinguishable from Haar-random, while the entanglement entropy is bounded away from volume-law. The latter property follows, as in previous examples, from the limit on the Schmidt rank of the state imposed by the tensor network structure. Namely, we have $S(A)\leq\min_{\gamma_{A}}|\gamma_{A}|\log(\chi)$, where $\gamma_{A}$ is a curve that separates $A$ from $\bar{A}$ in the network, and $|\gamma_{A}|$ is the number of bonds it cuts; see Fig. 3. Generally on hyperbolic tilings one has $\min_{\gamma_{A}}|\gamma_{A}|\sim\log|A|$, and thus sub-volme-law scaling, as claimed.
@@
A key property of holographic states (and random circuits or tensor networks more generally Nahum et al. (2017); Zhou and Nahum (2019); Li and Fisher (2021); Li et al. (2023)) is that their entropy approximately saturates the minimum-cut upper bound. Precisely, the Ryu-Takayanagi formula Ryu and Takayanagi (2006a, b) states that
@@
where $G$ is Newton’s constant in the bulk gravitational theory. We can prove that our ensemble $\mathcal{E}_{\rm holo}$ is indeed holographic in this sense. In Note (2) we prove a lower bound on the average entropy which, together with the aforementioned upper bound form the Schmidt rank, yields the following:
@@
For any subsystem $A$, the holographic pseudorandom tensor network ensemble $\mathcal{E}_{\rm holo}$ obeys
$\mathbb{E}_{\psi\sim\mathcal{E}_{\rm holo}}[S(A)]=|\gamma_{A}|\log(\chi)[1+o(1)]+O(N/\chi).$ (9)
Up to terms that vanish in the thermodynamic limit, this is a RT formula with “Newton’s constant” $G=4/\log(\chi)=o(1/\log(N))$.
Discussion.—We have constructed new models of pseudoentangled quantum state ensembles by combining pseudorandom unitaries and isometric tensor networks. Compared to prior constructions of pseudoentanglement, this approach provides more flexibility in the choice of entanglement structures that may be computationally hidden. These include pseudo-area-law scaling in any dimension as well as holographic entanglement which obeys a Ryu-Takayanagi formula, Eq. (9). The implications of our construction for holography remain to be explored, as well as possible extensions of our construction to more general and non-isometric holographic codes Akers et al. (2022); DeWolfe and Higginbotham (2023).
Given the prominent role of tensor networks in many-body physics (from exact representations of local Hamiltonian ground states to variational ansätze for time evolution to models of holography) this approach facilitates the integration of computational pseudorandomness in diverse physical settings. This paves the way for applications to the hardness of learning various properties of structured many-body states, such as Hamiltonian eigenstates or states prepared by specific types of non-thermalizing quantum dynamics. We leave these interesting directions to future work.
Note added.—While working on this manuscript we became aware of a forthcoming work by Akers, Bouland, Chen, Kohler, Metger and Vazirani Akers et al. (2024) that introduces a different construction of holographic pseudoentangled states. Additionally, during or after completion of this manuscript, several related works appeared on arXiv: Ref. Engelhardt et al. (2024) by Engelhardt et al introduced a different construction of pseudoentangled holographic states; Ref. Schuster et al. (2024) by Schuster, Haferkamp and Huang introduced a construction of pseudoentangled states based on shallow circuits of PFC PRUs—their 1D construction is similar, but not identical, to our MPS construction, and the techniques used in our proofs are different; finally Ref. Lami et al. (2024) by Lami, De Nardis and Turkeshi analitically and numerically studied random MPS and PEPS states in the regime of $\chi={\rm poly}(N)$, finding results that become consistent with ours when extrapolated to $\chi=\omega({\rm poly}(N))$.

### v3

@@ Abstract
Introduction.—Entanglement is a fundamental resource in quantum information science and an important principle in theoretical physics, from condensed matter to gravity Horodecki and Horodecki (2009); Kitaev and Preskill (2006); Levin and Wen (2006); Swingle (2012); Maldacena and Susskind (2013); Liu and Suh (2014); Kaufman et al. (2016); Abanin et al. (2019). At the same time, the entanglement structure of large many-body states is very challenging to observe. Specifically, there exist $N$-qubit states with very low entanglement that are computationally indistinguishable (i.e., indistiguishable to any observer whose resources scale polynomially in $N$) from states with much higher entanglement, such as Haar-random states. This property is knowns as pseudoentanglement Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Jeronimo et al. (2024); Bouland et al. (2023).
@@
Motivated originally by ideas in quantum cryptography Ji et al. (2018), the concept of pseudoentanglement enables connections to many-body physics in contexts including thermal equilibrium in many-body dynamics Feng and Ippoliti (2024), learning of Hamiltonian ground states Bouland et al. (2023), and holography Aaronson et al. (2023), where it may have implications for the complexity of the AdS/CFT “dictionary” Bouland et al. (2019). In general, it is a powerful tool to demonstrate the existence of entanglement structures or other quantum features that are computationally inaccessible Gu et al. (2024a, b), and thus contributes to sharpen the limits of what is learnable in realistic quantum experiments.
@@
Existing constructions of pseudoentangled state ensembles are based on subset-phase-states, |$\ket{\psi_{\mathcal{S},f}}\propto\sum_{x\in\mathcal{S}}(-1)^{f(x)}\ket{x}$, where under suitable constraints on the Boolean function $f:\{0,1\}^{N}\to\{0,1\}$ and/or the subset $\mathcal{S}\subseteq\{0,1\}^{N}$ one can prove computational pseudorandomness and enforce an upper bound on entanglement either on specific cuts or on all cuts at once Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Jeronimo et al. (2024). These constructions are very powerful and in some cases achieve the theoretical maximum amount of hideable entanglement. However, they natively represent unstructured states where all $N$ qubits are on the same footing, without e.g. a notion of locality. This makes them well suited to emulate maximally random states such as those that describe infinite-temperature equilibrium Feng and Ippoliti (2024), but limits their ability to mimic more structured systems. While it is possible to tailor phase-states to reflect some locality structure Bouland et al. (2023), it remains not clear in general how to mimic physical systems of interest, such as holographic ones.
@@
Motivated by these considerations, in this work we introduce new families of pseudoentangled states with a highly flexible entanglement structure. We leverage tensor networks Cirac et al. (2021); Evenbly and Vidal (2011); Zaletel and Pollmann (2020); Jahn and Eisert (2021); Haghshenas et al. (2022), which allow us to hard-code locality and the associated entanglement (pseudo-)area-law into our states. Recently proposed efficient constructions of pseudorandom unitaries Metger et al. (2024) then allow for the efficient preparation of pseudoentangled states whose true (computationally hidden) entanglement structure is that of an arbitrary tensor network that admits an isometric realization Zaletel and Pollmann (2020); Haghshenas et al. (2022); Soejima et al. (2020); Tepaske and Luitz (2021). These include many examples of interest, from matrix product states (MPS) to projected entangled pair states (PEPS) in any dimension Cirac et al. (2021); Verstraete et al. (2006); Schuch et al. (2007) and even holographic tensor network states Pastawski et al. (2015); Hayden et al. (2016); Jahn and Eisert (2021); Vasseur et al. (2019). While random tensor networks have been extensively studied in various contexts Hayden et al. (2016); Brandao et al. (2016); Nahum et al. (2017); Qi (2013); Vasseur et al. (2019), establishing pseudoentanglement requires new technical results in a distinct scaling regime for the bond dimension and the number of state copies, which we address in this work.
@@
in terms of the ‘moment operators’ $\rho^{(m)}_{\mathcal{E}}=\mathbb{E}_{\psi\sim\mathcal{E}}[\outerproduct{\psi}{\psi}^{\otimes m}]$. This means that an observer with access to polynomially many copies $m$ of a given state would have no way of deciding which ensemble it came from in polynomial time. The constraint on entanglement is simply that, for some extensive subsystem $A$ and with high probability over the state $\psi\sim\mathcal{E}$, the entanglement entropy $S(A)$ is bounded above by a function $f(N)\leq o(N)$. By comparison, the scaling of entanglement entropy in Haar-random states (known as the Page curve Page (1993)) is $S(A)=\Theta(N)$ for any extensive subsystem. The discrepancy between true and apparent entanglement is called the pseudoentanglement gap. The theoretically maximal pseudoentanglement gap is $f(N)=\omega(\log N)$ vs $\Theta(N)$, and can be saturated by random subset-phase states Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Jeronimo et al. (2024); Bouland et al. (2023).
@@
Pseudoentangled matrix product states.—We first exhibit our construction on the simplest tensor network: the one-dimensional matrix product state (MPS). A MPS, sketched in Fig. 1(a), is a many-body state whose amplitudes are expressed as $\innerproduct{i_{1},\dots i_{N}}{\psi}={\rm Tr}(A^{[1]}_{i_{1}}\cdots A^{[N]}_{i_{N}})$, with each $A^{[j]}_{i_{j}}$ a $\chi\times\chi$ matrix; $\chi$ is a numerical cut-off known as the bond dimension, which upper-bounds the Schmidt rank of the state about any bond on the 1D line and thus imposes an area-law, Fig. 1(b). A standard route to prepare MPSs in the lab is to use ‘staircase’ circuits  [^2] This approach is also known as “holographic” Foss-Feig et al. (2021); Anand et al. (2023), though not to be confused with the other usage of the word in this paper., where one associates $A_{i,\alpha,\beta}^{[j]}=\bra{i,\beta}U^{[j]}\ket{0,\alpha}$, with $U^{[j]}$ a unitary gate acting on a qubit and a $\chi$-state qudit. This is sketched in Fig. 1(c). In this work we take $\chi=2^{\nu}$, so each $U^{[j]}$ is a $(\nu+1)$-qubit gate. The staircase circuit comprises $L$ gates; for the last gate, all $\nu+1$ output qubits become part of the physical system, which in all is made of $N=L+\nu$ qubits. We take our unitary gates $\{U^{[j]}\}_{j=1}^{L}$ to be drawn from the recently introduced ‘PFC ensemble’ Metger et al. (2024) of pseudorandom unitaries (PRUs). These are the composition of a pseudorandom permutation (P), a pseudorandom binary phase function (F), and a random Clifford unitary (C)  [^3] The sampling of random Clifford unitaries is efficient as shown in Ref. van den Berg (2021), which give rise to efficiently implementable transformations that are computationally indistinguishable from Haar-random unitaries (see  [^4] See Supplementary Information. The Supplementary Information contains Ref. Collins and Matsumoto (2017)., Metger et al. (2024) for more details). We denote the resulting ensemble of $N$-qubit pure states by $\mathcal{E}_{\rm pPFCMPS}$ (pseudorandom PFC MPS).
@@
If $\omega({\rm poly}(N))\leq\chi\leq 2^{o(N)}$, the pseudorandom MPS ensemble $\mathcal{E}_{\rm pPFCMPS}$ is pseudoentangled.
@@
We split this task in three parts. (I) We remove the computational derandomization, which was only needed to enable efficient preparation. By definition of the PFC ensemble Metger et al. (2024), the pseudorandom ensemble $\mathcal{E}_{\rm pPFCMPS}$ is computationally indistinguishable from its random counterpart $\mathcal{E}_{\rm rPFCMPS}$, where the permutation and phase functions in the PFC gates are taken to be truly random instead of pseudorandom. (II) We prove that $\mathcal{E}_{\rm rPFCMPS}$ is indistinguishable from its counterpart made of Haar-random unitaries, which we term $\mathcal{E}_{\rm rMPS}$:
@@
We have $\|\rho^{(m)}_{\mathcal{E}_{\rm rPFCMPS}}-\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}\|_{\rm tr}\leq O(Nm/\sqrt{\chi})$.
The proof of this fact follows easily from the properties of the PFC ensemble, and is reported in Note (4). (III) We finally show that the random MPS ensemble $\mathcal{E}_{\rm rMPS}$ is indistinguishable from the ensemble of global Haar-random states. This is the main technical contribution of our work:
@@
    the $m$-th twirling channel for the Haar measure is given by $\Phi^{(m)}_{\rm Haar}(O)=\mathbb{E}_{U\sim{\rm Haar}}[U^{\otimes m}O(U^{\dagger})^{\otimes m}]=\sum_{\sigma,\tau\in S_{m}}\textsf{Wg}_{q}(\sigma\tau^{-1}){\rm Tr}(\hat{\tau}^{\dagger}O)\hat{\sigma}$, where $\textsf{Wg}_{q}(\sigma\tau^{-1})$ is the Weingarten function, see Note (4).
@@
where $D=2^{N}=2^{L}\chi$ is the physical Hilbert space dimension, each $\hat{\sigma}$ denotes a replica permutation for a single qubit ($i<L$) or $2\chi$-state qudit ($i=L$), and $e^{-E[\{\sigma_{i},\tau_{i}\}]}$ is a “Boltzmann weight” (not necessarily positive) of a configuration of $S_{m}$-valued spins on a line, sketched in Fig. 1(d). This way of writing the operator is reminiscent of a partition function Hunter-Jones (2019) and will be helpful in our analysis.
@@
We first show that $\Delta_{m}^{\rm u.}$ is small. The Boltzmann weight of the uniform spin configuration is $e^{-E[\{\sigma,\sigma\}]}=\chi^{m(L-1)}\textsf{Wg}_{2\chi}(e)^{L}$ (independent of $\sigma$), with $e$ the identity permutation. We show in Note (4) that $\textsf{Wg}_{2\chi}(e)=(2\chi)^{-m}[1+O(m^{2}/\chi)]$; therefore $e^{-E[\{\sigma,\sigma\}]}=D^{-m}[1+O(m^{2}/\chi)]$. At the same time, by the ‘birthday asymptotics’ we have $f_{D,m}=D^{-m}[1+O(m^{2}/\chi)]$, so overall
@@
Holographic tensor networks.—An important application of our construction is to tensor network models of holography, where a quantum system is viewed as the “boundary” of a higher-dimensional “bulk” space, and its entanglement structure reflects the geometry of the bulk. The existence of holographic pseudoentangled states, capable of hiding the bulk geometry from observers in the boundary quantum system, may have implications for the complexity of the holographic “dictionary” (the mapping between observables in the bulk and boundary theories) Bouland et al. (2019); Aaronson et al. (2023). Indeed, the dictionary is conjectured Susskind (2016) to relate an observable that is easy to measure (wormhole volume in the bulk gravity theory) to one that is extremely hard to measure (circuit complexity in the boundary quantum theory), seemingly violating standard complexity-theoretic assumptions. One proposed resolution is that the dictionary itself may be highly complex Bouland et al. (2019). There are two stages where this complexity might arise for an observer in the boundary theory: learning the boundary entanglement structure, and using it to reconstruct the bulk geometry. A holographic pseudoentangled state would ensure hardness of the former step, potentially allowing the latter step to be efficient (as was recently argued in a toy model Aaronson and Pollack (2023)). For this reason, Ref. Aaronson et al. (2023) asked whether it is possible to construct holographic pseudoentanglement.
@@
Holographic tensor networks have emerged as useful toy models of the AdS/CFT correspondence and its links to quantum error correction Pastawski et al. (2015); Swingle (2012); Qi (2013); Hayden et al. (2016); Qi et al. (2017); Jahn and Eisert (2021). They generally feature tensors arranged on a tiling of $2D$ hyperbolic space terminating with “dangling” legs on the boundary which represents the physical system of interest. This structure is very well suited to our construction, since it encodes the bulk isometrically into the boundary; i.e., the tensor network may be viewed as a circuit propagating information unitarily outward from the center of the hyperbolic space toward its boundary.
@@
Here we focus for concreteness on the $\{6,4\}$ tiling of the hyperbolic plane with 6-leg tensors (3-qudit unitary gates), without bulk degrees of freedom. This is sketched in Fig. 2(b). Placing a PRU of dimension $\chi^{3}$ on each hexagon defines a state ensemble $\mathcal{E}_{\rm holoTNS}$. Our general proof extends to this ensemble and is sufficient to show pseudoentanglement: $\mathcal{E}_{\rm holoTNS}$ is computationally indistinguishable from Haar-random, while the entanglement entropy is bounded away from volume-law. The latter property follows, as in previous examples, from the limit on the Schmidt rank of the state imposed by the tensor network structure. Namely, we have $S(A)\leq\min_{\gamma_{A}}|\gamma_{A}|\log(\chi)$, where $\gamma_{A}$ is a curve that separates $A$ from $\bar{A}$ in the network, and $|\gamma_{A}|$ is the number of bonds it cuts; see Fig. 3. Generally on hyperbolic tilings one has $\min_{\gamma_{A}}|\gamma_{A}|\sim\log|A|$ for contiguous $A$, and thus sub-volume-law scaling, as claimed.
@@
However this does not yet establish holographic entanglement, as it is only an upper bound. A key property of holographic states (and random circuits or tensor networks more generally Nahum et al. (2017); Zhou and Nahum (2019); Li and Fisher (2021); Li et al. (2023)) is that their entropy approximately saturates this minimum-cut upper bound. Precisely, the Ryu-Takayanagi formula Ryu and Takayanagi (2006a, b) states that
@@
where $G$ is Newton’s constant in the bulk gravitational theory. We can prove that our ensemble $\mathcal{E}_{\rm holoTNS}$ is indeed holographic in this sense. In Note (4) we prove a lower bound on the average entropy which, together with the aforementioned upper bound form the Schmidt rank, yields the following:
@@
For any subsystem $A$, the holographic pseudorandom tensor network ensemble $\mathcal{E}_{\rm holoTNS}$ obeys
$\mathbb{E}_{\psi\sim\mathcal{E}_{\rm holoTNS}}[S(A)]=|\gamma_{A}|\log(\chi)[1+o(1)]+O(N/\chi).$ (9)
Up to terms that vanish in the thermodynamic limit, this is a RT formula with “Newton’s constant” $G=4/\log(\chi)=o(1/\log(N))$. Our proof follows the standard approach for holographic random tensor networks Hayden et al. (2016); the reduction of our PRU-based construction to a random calculation is made possible by the fact that the PFC ensemble forms an exact 2-design (due to the random Clifford part).
Combining Theorem 2 with the proven computational indistinguishability Note (4), we obtain the existence of holographic pseudoentangled states, answering the question posed in Ref. Aaronson et al. (2023).
Discussion.—We have constructed new models of pseudoentangled quantum state ensembles by combining pseudorandom unitaries and isometric tensor networks. Compared to prior constructions of pseudoentanglement, this approach provides more flexibility in the choice of entanglement structures that may be computationally hidden. These include pseudo-area-law scaling in any dimension as well as holographic entanglement which obeys a Ryu-Takayanagi formula, Eq. (9). It would be interesting to extend our construction to more general and non-isometric holographic codes Akers et al. (2022); DeWolfe and Higginbotham (2023). More broadly, the implications of our construction for holography remain to be explored; it is worth emphasizing that tensor networks are a very simplistic toy model of holography and further work is needed to extend these results toward more realistic models.
Given the prominent role of tensor networks in many-body physics (from exact representations of local Hamiltonian ground states to variational ansätze for time evolution to models of holography) our results facilitate the integration of computational pseudorandomness in diverse physical settings. This paves the way for applications to the hardness of learning various properties of structured many-body states, such as Hamiltonian eigenstates or states prepared by specific types of non-thermalizing quantum dynamics. We leave these interesting directions to future work.
Note added.—While working on this manuscript we became aware of related work by Akers, Bouland, Chen, Kohler, Metger and Vazirani Akers et al. (2024) that introduces a different construction of holographic pseudoentangled states. Additionally, during or after completion of this manuscript, several related works appeared on arXiv: Ref. Engelhardt et al. (2024) by Engelhardt et al introduced a different construction of pseudoentangled holographic states; Ref. Schuster et al. (2024) by Schuster, Haferkamp and Huang introduced a construction of pseudoentangled states based on shallow circuits of PFC PRUs—their 1D construction is similar, but not identical, to our MPS construction, and the techniques used in our proofs are different; finally Ref. Lami et al. (2024) by Lami, De Nardis and Turkeshi analitically and numerically studied random MPS and PEPS states in the regime of $\chi={\rm poly}(N)$, finding results that become consistent with ours when extrapolated to $\chi=\omega({\rm poly}(N))$.
Acknowledgements.